OpenStudy (anonymous):

what is the point slope equation of the line that has a slope of 1/8 and goes through the point (2,-3)?

OpenStudy (lgbasallote):

POINT SLOPE FORM \[\large y - y_1= m (x - x_1)\]

OpenStudy (anonymous):

is it y+3=1/8(x-2)?

OpenStudy (anonymous):

im not sure :l